OpenStudy (anonymous):

How would I solve |2x-4|=5?

OpenStudy (anonymous):

Make 2 seperate equtions and solve 1. 2x-4=5 2. 2x-4=-5

OpenStudy (john_es):

You can split in two equations. First you need to know where 2x-4 change its sign. \[2x-4=5\Rightarrow x=9/2\]Now you have to solve simultaneously two equations, \[x\geq9/2\Rightarrow2x-4=5\]\[x<9/2\Rightarrow -(2x-4)=5\]Remember to check that the solutions lies in the cited interval.

OpenStudy (anonymous):

Thanks! :D